OpenStudy (anonymous):

A boy throws a ball up into the air with a speed of 8.2 m/s. The ball has a mass of 0.3 kg. How much gravitational potential energy will the ball have at the top of its flight? (Assume there is no air resistance.) A. 19.3 J B. 10.1 J C. 8.7 J D. 15.2 J

OpenStudy (anonymous):

B. 10.1 J kinetic energy and potential energy are nonconvertible. When the ball is thrown it possess kinetic energy and when it reaches at the top the whole K.E. gets converted into P.E. Therefore P.E. i.e. K.E. \[=\frac{ 1 }{ 2 }mv ^{2}\] \[=\frac{ 1 }{ 2 }\times0.3\times(8.2)^{2}\]\[=10.086 \approx 10.1J\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Is that a correct answer? Please confirm I just wrote down what I remember from high-school.

OpenStudy (anonymous):

That was the correct answer!